OpenStudy (anonymous):

Solve the quadratic equation by completing the square. x^2-18x+67=0

OpenStudy (freckles):

\[x^2-18x=-67 \\ x^2-18x+? =-67 +?\] we need to add something on both sides such that we can write the left hand side as something square

OpenStudy (freckles):

\[x^2+kx+(\frac{k}{2})^2=(x+\frac{k}{2})^2\]

OpenStudy (freckles):

notice the number in front of x in your equation

OpenStudy (freckles):

that is your k

OpenStudy (anonymous):

Can I show you what I have so far?

OpenStudy (freckles):

yes totally

OpenStudy (anonymous):